OpenStudy (anonymous):

(I had my surgery done two days ago nd I'm still very slow, I will give a medal) There is a Saturday performance at a local theater. A theater charges $9 for an adult ticket and $4 for a child ticket. All 1020 tickets were sold and $7680 of revenue was brought in. How many adult tickets were sold?

OpenStudy (driftracer305):

do i give straight answer or u want explaination?

OpenStudy (joannablackwelder):

I think you will need to set up a system of equations. Do you have an idea of how to start?

OpenStudy (anonymous):

When you first look at the question, you should think, i have A (adult tickets) and C(children tickets). So you have *TWO UNKNOWNS*.... therefore you need..... (have a guess!)

OpenStudy (anonymous):

@Keridavis

OpenStudy (anonymous):

explanation and answer please. i can't focus well. meds are kicking in.

OpenStudy (anonymous):

wht ya have surgurey for and on?

OpenStudy (anonymous):

Tonsils and Adenoids

OpenStudy (aripotta):

ok so $9 for adults, $4 for children. $7680 in revenue. set up an equation equaling 7680 with 2 variables, a for adults and c for children. and then another equation equaling 1020 because that's how many tickets were sold total 9a + 4c = 7680 a + c = 1020 can you solve it from there?

OpenStudy (anonymous):

not really

OpenStudy (aripotta):

can you take the second equation and solve for c?

OpenStudy (anonymous):

hope you feel better

OpenStudy (aripotta):

9a + 4c = 7680 a + c = 1020 solve for c in the second equation first since c doesn't have a coefficient in that equation, which makes it easier for us. c = 1020 - a then plug that into the first equation 9a + 4(1020 - a) = 7680 distribute the 4 9a + 4080 - 4a = 7680 combine like terms 5a + 4080 = 7680 then work to get a by itself. subtract 4080 from both sides 5a = 3600 divide both sides by 5 a = 720

OpenStudy (anonymous):

Thank you so much.

OpenStudy (aripotta):

no problem, i hope you understand better :) oh and get well soon!

OpenStudy (anonymous):

Thank you so much and I'm trying. I'm just in alot of pain

OpenStudy (aripotta):

not to be a jerk or anything, but can you give me a medal please? :)

OpenStudy (aripotta):

i understand

OpenStudy (anonymous):

Of course.

OpenStudy (anonymous):

I hope you get well and stop hurting

OpenStudy (anonymous):

Thank you. I'm trying

OpenStudy (anonymous):

I know the feeling I had knee surgery 2 months ago